Arna is twice as old as Mae. In 4 years time, she will be five-thirds as old as Mae. How old is each now

with explanation pls
1
Expert's answer
2019-09-23T06:12:13-0400

Let Mae's current age be x years.

Since Arna is twice as old as Mae, Arna's present age is 2x years.


After 4 years, Mae's age will be x+4 years and Arna's age will be 2x+4 years.


As per given information, after 4 years, Arna will be five-thirds as old as Mae.


Thus, the equation is:


"2x + 4 = \frac{5}{3}*(x+4)"

"\therefore 3 * (2x+4) = 5*(x+4)"


"\therefore 6x + 12 = 5x + 20"


"\therefore 6x - 5x = 20 - 12"


"\therefore x = 8"


Hence 2x = 2 * 8 = 16 years


Thus, Mae's current age is 8 years and Arna's current age is 16 years.